# Profiling Bias in LLMs: Stereotype Dimensions in Contextual Word Embeddings

### v1

@@ 4.2 Gender Stereotype Profiles
For all twelve studied LLMs, we find statistically significant bias for gender-associated names, as evident in the 2D profiles for warmth and competence in Figure 3. In line with human bias found in studies of the stereotype content model (Fiske et al., 2002), the models highly agree on the relative associations of female names with warmth and male names with competence when using the gender-neutral generated contexts. GPTNeo poses an exception, where both dimensions are associated with female names, and for OlMo, the difference in competence is insignificant.
For the much smaller ‘vocabulary populations’ of gendered terms (nine terms per gender), warmth is the more relevant dimension than competence, with only the former being significant for all four Llama-3 models. For some models, e.g. GPT2, the gendered term differences are small, but the bias direction is very consistent when comparing name and term profiles. In Figure 4, we additionally see the projections of five individual terms beyond binary gender. Nonbinary and transgender-related terms are associated with lower warmth than binary term means, which was found for all newer models (studied variants of LLama3, Gemma, OLMo). For competence, there is no respective trend.
Zooming into the 7-dimensional stereotype profiles for Lama-3-8B (see Figure 1) and Llama-3.2-3B-instruct in Figure 5, we can perceive the previously shown associations in more detail. The profiles between Llama-3 and the newer and instruction-tuned 3.2 version are quite similar: Sociability and morality (warmth) are linked with female names, which is true for 10 and 11 of all studied models. Ability and agency (competence) are significantly related with male names, which is true for 9 and 7 of the studied models. Also for status, six models show a significant association with male names. Furthermore six models find male names to be more traditional on the political dimension, while there is no perceivable trend for religion. For gendered terms, there is only one clearly biased dimension, where seven models agree on a stereotypical association of sociability with female terms.
@@ 4.3 Context Examples and Bias across Layers
Comparing contexts across layers in Figure 6 we find that their influence on prediction accuracy highly depends on the model. For Llama-3-8B, the accuracy is quite stable compared to the smaller Gemma-2B variant, where we see much variation across layers and context types. For GPT2 the Reddit examples lead to much lower accuracy, while for BERT the no context condition performs markedly worse.
On the right in Figure 6 we see that bias across layers is rather consistent for all models, where higher (lower) values signify female (male) bias. Here, shown by generated examples, stereotypical associations permeate throughout the networks, also for Gemma, despite the unstable semantics indicated by the prediction task. In some cases, the first and last layers behave differently with a reversed direction of bias compared to the overall model.
Figure 6: Accuracy for warmth and competence across layers for different context examples (left) and gender bias across layers (right) with generated gender-neutral examples. Higher values signify bias towards female associated names/terms, lower values signify bias towards male associated names/terms.
@@ 5 Discussion
Our results provide substantial evidence of stereotype dimensions in the embedding space of LLMs and a gender bias that predominantly corresponds to the human bias found in studies of the stereotype content model (Fiske et al., 2002). For all models, female names are significantly associated with warmth, and for most models, male names are also associated with competence. There is less evidence for bias for the studied gendered terms, which in part is certainly due to the small groups of only nine terms per gender. The direction of gender differences is, however, overwhelmingly consistent.
@@
The analysis of contextual embeddings based on the stereotype content model can deliver the most robust results when analyzing large vocabulary groups and broader concepts. Statistical analysis is viable with the larger collection of names, compared to the small groups of binary gendered terms or the even shorter list of terms for nonbinary gender. While both levels of stereotype dimensions provide interesting results, a significant gender bias is most evident in the warmth and competence dimensions. These benefit from the larger numbers of pole words in the dictionaries, increasing the robustness of the concept representations. Furthermore, associations may generally be clearer when relating to broader concepts.
@@
Significant gender bias, however, may also occur on a more granular level. Other dimensions can be relevant depending on domains and tasks, such as progressive-traditional in the realm of politics, and the mode of projection can be chosen accordingly. A combined projection with other dimensions such as valence (unpleasantness vs pleasantness), as done by Omrani Sabbaghi et al. (2023), could provide even further insights.
@@
As term context can have a considerable influence on the stereotype dimensions, examples for pole and projected terms should be chosen deliberately. Gender-neutral context is a good standard choice, but measurement might best be conducted with domain-specific data when a specific use case exists. While we use neutral templates as contexts for the gendered terms and names, these could as well be sampled from a target domain or be generated to test specific scenarios. By introducing success in a historically male-dominated field to the term/name context, it could be tested if this leads to a penalty for females, as found in psychological studies (Heilman et al., 2004).
@@
While embedding-based methods for bias measurement have been critiqued for their remoteness from downstream applications (Gallegos et al., 2024) and are certainly no substitute for task-specific investigations, they have multiple advantages. First, the methodology does not depend on natural language datasets that can be leaked into training data and are therefore applicable to older and newer models alike. Second, the same stereotype dimensions can easily be used for bias mitigation (Ungless et al., 2022; Omrani et al., 2023), alleviating representational harm and the risk that this influences downstream behavior. Finally, our paper shows they can be exploited for intuitive visualizations exposing gender bias.
@@ 6 Conclusion
In this work, we profile twelve LLMs by means of the stereotype content model of social psychology (Fiske et al., 2002), thereby theoretically grounding the analysis, which has in the past been described as the missing link for bias measurements (Blodgett et al., 2020). We employ a polar transformation on the contextual embeddings, finding significant bias for gender-associated names and gendered terms along the stereotype dimensions, widely aligned with human bias.
Bias profiles based on the stereotype content model employ an intuitive scoring along scales of opposing concepts (e.g. low vs. high warmth), as proven effective by semantic differentials in human surveys (Osgood et al., 1957). The result is a highly visual solution for communicating bias to wider audiences and users of artificial intelligence.
@@ 7 Limitations
The bias profiles presented in this paper concern gender, but there is a whole range of biases to be profiled in LLMs to assess and communicate representational harms. The scope of analysis was also constrained to English contextual embedding spaces and needs to be extended to a multi-lingual setting in the future.
Furthermore, the focus of this paper was binary gender, with gender-associated names and gendered terms. While we projected a few terms for transgender and nonbinary gender to the stereotype dimensions, future analysis needs to extend the methodology for these smaller and diverse ’vocabulary populations’. Increasing the number of context samples for terms could have the potential for greater robustness and additional statistical tests.
Finally, no general bias measurement benchmark or method, including the one presented in this paper, precludes the absolute necessity of task-specific bias measurements. However, they can be a piece of the puzzle by revealing learned general bias tendencies and providing a means to mitigate and communicate them effectively.

### v2

@@ 4.2 Gender Stereotype Profiles
For all twelve studied LLMs, we find statistically significant bias for gender-associated names, as evident in the 2D profiles for warmth and competence in Figure 3. In line with human bias found in studies of the stereotype content model (Fiske et al., 2002), the models highly agree on the relative associations of female names with warmth and male names with competence when using the gender non-specific generated contexts. GPTNeo poses an exception, where both dimensions are associated with female names, and for OlMo, the difference in competence is insignificant.
For the much smaller ‘vocabulary populations’ of gendered terms (nine terms per gender), warmth is the more relevant dimension than competence, with only the former being significantly biased for all four Llama-3 models. For some models, e.g. GPT2, the gendered term differences are small, but the bias direction is very consistent when comparing name and term stereotype profiles. In Figure 4, we additionally see the projections of five individual terms beyond binary gender. Nonbinary and transgender-related terms are associated with lower warmth than binary term means, which was found for all newer models (studied variants of LLama3, Gemma, OLMo). For competence, there was no observable trend. For further and statistical analyses of individual terms and small vocabulary populations, future work needs to extend the context examples (as discussed in section 5).
Zooming into the 7-dimensional stereotype profiles for Lama-3-8B (see Figure 1) and Llama-3.2-3B-instruct in Figure 5, we can perceive the previously shown gender associations in more detail. The profiles between Llama-3 and the newer and instruction-tuned 3.2 version are quite similar: Sociability and morality (warmth) are linked with female names, which is true for 10 and 11 of all studied models. Ability and agency (competence) are significantly related with male names, which is true for 9 and 7 of the studied models. Also for status, six models show a significant association with male names. Furthermore six models find male names to be more traditional on the political dimension, while there is no perceivable trend for religion. For the small populations of binary gendered terms, there is only one clearly biased dimension, where seven models agree on a stereotypical association of sociability with female terms.
Figure 4: 2D Stereotype profile for Llama-3-8B (see Figure 3) with additional projections of individual nonbinary terms. LW/HW = Low/High Warmth. LC/HC = Low/High Competence.
Figure 5: 7D stereotype profiles for 100 female/male-associated names (left) and 9 female/male gendered terms (right) for Llama-3.2-3B-instruct. *Statistically significant differences (p<0.05).
(a) Accuracy for the direction prediction task across layers, with different context examples. Additional terms are projected to the stereotype dimensions; positive/negative values predict high/low direction.
(b) Gender bias across layers with generated gender non-specific examples. Higher values signify bias towards female-associated names/terms; lower values signify bias towards male-associated names/terms.
Figure 6: Layerwise visualization of prediction accuracy and gender bias for selected models.
@@ 4.3 Context Examples and Bias across Layers
Comparing the performance for generated, dictionary, Reddit examples, and no context across layers in 6(a), we find that the influence of context on the new term prediction task depends on the model. For Llama-3-8B, the accuracy seems quite stable compared to the smaller Gemma-2B variant, where we see high variation across layers and context types. However, as elaborated in subsection 4.1, while the cut-off value of zero works well for most, including all Llama models, Gemma could better discriminate between low/high labeled terms by a different cut-off value. For GPT2, the Reddit examples lead to much lower accuracy, while for BERT, the no context condition performs markedly worse. Overall the concepts of warmth and competence behave similarly throughout the layers.
On the right in 6(b), we see that bias across layers is rather consistent for all models, where higher values signify bias towards female-associated names/terms and lower values signify bias towards male-associated names/terms. Shown by the example of the generated contexts, stereotypical associations permeate throughout the networks. In some cases, the first and last layers behave differently, with a reversed bias direction compared to the overall model.
@@ 5 Discussion
Our results provide substantial evidence of stereotype dimensions in the embedding space of LLMs and a gender bias that predominantly corresponds to the human bias found in studies of the stereotype content model (Fiske et al., 2002). For all studied models, female names are relatively associated with higher warmth, and for most models, male names are associated with higher competence. There is less evidence of bias for the studied gendered terms, which in part is likely due to the small groups of only nine terms per gender. The direction of gender differences is, however, overwhelmingly consistent. We furthermore find stereotype dimensions and bias across layers, in line with prior work that semantics are spread throughout the network (Tenney et al., 2019a).
The projection of contextual embeddings based on the stereotype content model can deliver robust insights when analyzing large vocabulary groups. As the magnitude of the values depends on the properties of the original embedding space, statistical analysis is employed to assert the significance of gender differences. This is viable with the larger collection of gender-associated names, also providing context for the differences between binary gendered terms. For the analysis of small vocabulary groups or individual terms, e.g. for comparing terms of binary and nonbinary gender, increasing the number of context examples offers potential for statistical tests.
@@
While both 2D and 7D stereotype dimensions provide interesting results, a significant gender bias is most evident in the warmth and competence dimensions. These benefit from the larger numbers of low and high-rated words in the dictionaries, increasing the robustness of the concept representations. Likely associations are also more stable when relating to broader concepts. Therefore, the high-level projection is a suitable first level of analysis and starting point for bias mitigation.
@@
Significant gender bias, however, may also occur on a more granular level. Different dimensions can be relevant depending on domains and tasks, such as progressive-traditional in the realm of politics, and the mode of projection can be easily adapted with the presented methodology. A combined projection with other dimensions such as valence (unpleasantness vs. pleasantness) (see e.g. Omrani Sabbaghi et al. (2023)), could provide even further insights.
@@
As we have shown, the term context can have a considerable influence on the behavior of the stereotype dimensions. Thus, examples for pole and projected terms should be chosen deliberately. Gender non-specific context is our default choice because no additional bias is introduced through the examples and we get a clearer picture of the bias already present within the pre-trained embeddings. Even smaller open-source LLMs are now able to provide examples with this property at scale. However, measurement is certainly best conducted with domain-specific data when a specific use case exists. While we use simple templates as contexts for the gendered terms and names, these could as well be sampled from a target domain or be generated to test specific scenarios. For example, similar to May et al. (2019), this could involve introducing success in a historically male-dominated field to the term/name context, to test if a penalty exists for females, as found in psychological studies (Heilman et al., 2004).
@@
While embedding-based methods for bias measurement have been critiqued for their remoteness from downstream applications (Gallegos et al., 2024), and are certainly no substitute for task-specific investigations, they have multiple advantages. First, the methodology does not depend on natural language datasets that can be leaked into training data and are therefore applicable to older and newer models alike. Second, the same stereotype dimensions can easily be used for bias mitigation (Ungless et al., 2022; Omrani et al., 2023), alleviating representational harm and the risk that it influences downstream behavior. Finally, our paper shows they can be exploited for intuitive visualizations exposing gender bias.
@@ 6 Conclusion
In this work, we profile gender stereotypes in twelve LLMs by means of the stereotype content model of social psychology (Fiske et al., 2002), thereby theoretically grounding the analysis, which has in the past been described as the missing link for bias measurements (Blodgett et al., 2020). By a matrix transformation, the opaque contextual embeddings of the models reveal interpretable stereotype dimensions. Along the two major dimensions of warmth and competence, we find significant bias for gender-associated names and some evidence of bias for gendered terms, widely aligned with stereotypes found in human studies.
The shown presence of stereotype dimensions in LLMs is a comprehensible replication of semantics in human language, however, the differential associations of social groups along these dimensions constitute a representational harm. Equal treatment starts with equal representation; stereotypes already statistically significant in embedding space come with the risk of being exploited in downstream tasks, which could lead to different and unfair treatment of social groups. While the first access point would be the training data itself, the embedding space allows a quantification of patterns that is useful for bias assessment and mitigation. The analysis of embedding spaces by interpretable dimensions provides a means to evaluate both functional semantics and harmful associations that should be ‘unlearned’ to prevent their propagation.
Awareness of bias in LLMs needs to be increased beyond the expert audience, as biased models are already deployed, and completely debiased models may hardly be attainable, as they are trained on vast amounts of biased human-created data. The here presented bias profiles based on the stereotype content model employ an intuitive scoring along meaningful scales of opposing concepts (e.g. low vs. high warmth), as proven effective by semantic differentials in human surveys (Osgood et al., 1957). The result is a highly visual solution for communicating bias to wider audiences and users of artificial intelligence.
@@ 7 Limitations
The bias profiles presented in this paper concern only gender, but there is a whole range of biases to be profiled in LLMs to evaluate and communicate representational harms. The scope of analysis was also constrained to English contextual embedding spaces and needs to be extended to a multi-lingual setting in the future.
Furthermore, the focus of this paper was binary gender, with historically gender-associated names and gendered terms. While we projected a few terms for transgender and nonbinary gender to the stereotype dimensions, future analysis needs to extend the methodology for these smaller and diverse ‘vocabulary populations’. Increasing the number of context samples for terms offers potential for greater robustness and applicability of statistical tests.
Finally, no general bias measurement benchmark or method, including the one presented in this paper, precludes the absolute necessity of task-specific bias measurements. However, they can be a piece of the puzzle by revealing learned general bias tendencies and providing a means to mitigate and communicate these effectively.